The 3rd and 4th terns of a geometric progression are 12 and 6 respectively. What are the first five terms?
Answer by solve_for_x(190)

The common factor of the progression is 6/12 = 1/2

Since the third term is 12, the second term is 12/(1/2) = 24.

Since the second term is 24, the first term is 24/(1/2) = 48.

And since the 4th term is 6, the fifth term is 6(1/2) = 3.

The first five terms of the progression are then:

48, 24, 12, 6, 3